# Density Descent for Diversity Optimization

### v1

@@ 4 Density Descent Search
The second variant, DDS-CNF, estimates the density in feature space by learning a CNF between the standard normal distribution $\mathcal{N}(\bm{0},{\bm{I}})$ and the observed feature space distribution. Similar to DDS-KDE, the feature space distribution is represented by the buffer ${\mathcal{B}}$. We compute a density estimate at any location in the feature space by integrating an ordinary differential equation (Lipman et al.,, 2023). Applying techniques from prior work (Lipman et al.,, 2023), we represent the CNF with a neural network, and on each iteration, we finetune the network on the new distribution of features contained in the buffer.
KDE and CNF differ in how easily their smoothness can be controlled. On one hand, the shape of the KDE can be controlled with the bandwidth hyperparameter. Higher bandwidth leads to a smoother density estimate but conceals modes of the true density function, as illustrated in 1(a). On the other hand, while CNF does not require selecting a bandwidth hyperparameter, the smoothness of the density estimation cannot be easily controlled, which undermines the performance of the algorithm (Sec. 6.3). We discuss the impacts that the smoothness of the density estimate has on DDS in Sec. 5.
Feature Buffer. To provide the basis for density estimates in feature space, DDS maintains a buffer ${\mathcal{B}}$ that stores the features of sampled solutions (Algorithm 1). This buffer represents the observed distribution of the feature space and is updated every time new solutions and features are discovered. In theory, the buffer can have infinite capacity, storing every feature ever encountered. In practice, the buffer can only retain a finite number of features due to computation and memory limitations. To decide which features to retain in the buffer, we manage the buffer with an optimal reservoir sampling algorithm (Li,, 1994). This algorithm updates the buffer with online samples that accurately represent the distribution of features discovered so far by DDS (Algorithm 1).
Optimizer. DDS guides an adaptive optimizer to discover solutions in low-density regions of the feature space. Examples of such optimizers include xNES (Glasmachers et al.,, 2010) and Adam (Kingma and Ba,, 2015). We select CMA-ES as the underlying optimizer due to its reputation as a state-of-the-art optimizer (Hansen,, 2016) and its high performance in existing QD algorithms (Fontaine et al.,, 2020; Fontaine and Nikolaidis,, 2021, 2023).
Passive Archive. Following prior work (Pugh et al.,, 2015), to track how much of the feature space has been explored, DDS inserts all discovered solutions and features into a passive MAP-Elites-style (Sec. 3) archive $\mathcal{A}$. While DDS itself only uses the archive to store solutions and features, the archive is useful when computing metrics in experiments (Sec. 6.1).
@@
1 DDS (ϕ()bold-$\bm{\phi}(\cdot)$, $density$, $b,N,\lambda,{\bm{\mu}}_{0},\sigma$)
@@
2       Initialize CMA-ES search point ${\bm{\mu}}:={\bm{\mu}}_{0}$, search direction ${\bm{\Sigma}}=\sigma{\bm{I}}$, and internal parameters ${\bm{p}}$ Initialize empty archive $\mathcal{A}$ Initialize empty buffer ${\mathcal{B}}$ of size $b$ for $iter\leftarrow 1$ to $N$ do
@@
4                   ${\bm{\theta}}_{i}\sim\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$ ${\bm{\phi}}_{i}\leftarrow{\bm{\phi}}({\bm{\theta}}_{i})$ $D_{i}\leftarrow density({\bm{\phi}}_{i})$ Add $({\bm{\theta}}_{i},{\bm{\phi}}_{i})$ to archive $\mathcal{A}$
5             end for
6            Update buffer ${\mathcal{B}}$ with ${\bm{\phi}}_{1..\lambda}$ via reservoir sampling Update $density$ with the new buffer ${\mathcal{B}}$ Rank ${\bm{\theta}}_{i}$ in ascending order by $D_{i}$ Adapt CMA-ES parameters ${\bm{\mu}},{\bm{\Sigma}},{\bm{p}}$ based on density ranking $D_{i}$
7       end for
8
9
@@
5 Connection between Novelty Search and Kernel Density Estimates
@@
Stability Under Non-stationarity. DO algorithms such as NS, DDS, and CMA-MAE gradually learn a non-stationary density representation (e.g., KDE or histogram) as they explore the feature space. However, drastic changes in the representation present significant challenges for adaptive optimizers, such as Adam (Kingma and Ba,, 2015) and CMA-ES. To characterize the extent of change in the density estimate, we appeal to the notion of uniform stability (Hardt et al.,, 2016), defined as follows:
@@
Let a function $F({\bm{x}};{\mathcal{B}}):\mathbb{R}^{d}\to\mathbb{R}$ be parameterized by a set ${\mathcal{B}}\in\mathbb{R}^{d}$. We say that $F$ is $\epsilon$-uniformly stable if for all ${\mathcal{B}},{\mathcal{B}}^{\prime}\subseteq\mathbb{R}^{d}$, where ${\mathcal{B}}$ and ${\mathcal{B}}^{\prime}$ differ by at most one element, we have
$\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left|F({\bm{x}};{\mathcal{B}})-F({\bm{x}};{\mathcal{B}}^{\prime})\right|\leq\epsilon$ (3)
We note that uniform stability has close connections to influence functions in statistics, which measures how much an estimator deviates from the ground truth when given a subset of the data (Cook and Weisberg,, 1980). In this interpretation, uniform stability is an upper-bound of the (empirical) influence function.
We prove that KDE has uniform stability directly proportional to the size of its feature buffer and bandwidth:
@@
Higher bandwidth makes the function more stationary and thus more suited for adaptive optimizers. However, higher bandwidth also leads to over-smoothing of the density estimate, which conceals modes of the true density function (Chen,, 2017) (1(a)). Thus, selecting an optimal bandwidth for DDS requires a fine balance between the accuracy and uniform stability of the KDE.
@@
Therefore, for unbounded feature spaces, the uniform stability of novelty score is also unbounded, and for bounded feature spaces, novelty score has uniform stability directly proportional to $k$.
When the feature space is bounded, as in our experiments, KDE has a stronger stability guarantee than novelty score for bandwidth $h\geq\frac{k}{|{\mathcal{B}}|}$. Following this insight, we select a bandwidth satisfying this inequality with the bandwidth selection mechanism in Appendix B. Our theoretical results are corroborated by our experiments in Sec. 6, which demonstrate that DDS-KDE outperforms NS in all domains on all metrics.
@@
6 Experiments
We compare the performance of DDS-KDE and DDS-CNF with the QD algorithms MAP-Elites (line), CMA-ME, and CMA-MAE, and with NS using CMA-ES as the underlying optimizer. MAP-Elites (line) describes the implementation of MAP-Elites with the Iso+LineDD operator (Vassiliades and Mouret,, 2018). All algorithms are implemented with pyribs (Tjanaka et al.,, 2023).
Our experiments include canonical benchmark domains from QD and DO: Linear Projection (Fontaine et al.,, 2020), Arm Repertoire (Vassiliades and Mouret,, 2018), and Deceptive Maze (Lehman and Stanley,, 2011). As discussed in Sec. 2, we convert QD domains into DO domains by setting the objective to be constant, effectively removing the importance of solution quality. For Deceptive Maze, we use the implementation in Kheperax (Grillotti and Cully,, 2023). Furthermore, to address the challenges posed by high-dimensional feature spaces, we introduce a new domain, Multi-feature Linear Projection, that generalizes Linear Projection to high-dimensional feature spaces.
@@ 6.1 Experimental Design
Dependent Variables. We compute the archive coverage as the number of occupied cells in the archive divided by the total number of cells. To compare the coverage of tessellation-based algorithms (CMA-MAE, CMA-ME, and MAP-Elites) with that of non-tessellation-based algorithms (DDS-KDE, DDS-CNF, and NS), we track the coverage of all algorithms on a passive archive $\mathcal{A}$ tessellated by a $100\times 100$ grid. For the high-dimensional feature space experiment, we track coverage with a centroidal Voronoi tessellation with 10,000 cells (Vassiliades et al.,, 2018).
Following prior work (Gomes et al.,, 2015), we also assess the ability of each algorithm to uniformly explore the feature space. Thus, we measure the cross-entropy between a uniform distribution and the distribution of sampled features. Let $N_{e}$ be the total times throughout the entire search that solutions were discovered in cell $e\in\{1,\dots,l\}$ in the passive archive, and $N_{\mathrm{total}}=\sum_{e=1}^{l}N_{e}$. The cross-entropy score is defined as:
$\mathrm{CE}=-\sum_{e=1}^{l}\frac{1}{l}\log\left(\frac{N_{e}}{N_{\mathrm{total}}}\right)$ (4)
$CE$ achieves its minimum value when $N(\cdot)$ is uniformly distributed. For all passive archives in our experiments, this minimum is 9.21 for $l=$10,000$$ cells.
6.2 Domain Details
Linear Projection (LP). LP is a QD domain where ${\bm{\phi}}$ induces high distortion by mapping an $n$-dimensional solution space to a 2D feature space (Fontaine et al.,, 2020). A harsh penalty is applied outside the bounds of the feature space to hinder exploration near the bounds. The feature function ${\bm{\phi}}:\mathbb{R}^{n}\to\left[-5.12\cdot\frac{n}{2},5.12\cdot\frac{n}{2}\right]^{2}$ is defined as:
${\bm{\phi}}({\bm{\theta}})$ $=\left(\sum_{i=1}^{\frac{n}{2}}\mathrm{clip}(\theta_{i}),\sum_{i=\frac{n}{2}+1}^{n}\mathrm{clip}(\theta_{i})\right)$ (5)
$\mathrm{clip}(\theta_{i})$ =$=\begin{cases}\theta_{i}&\text{if}\;|\theta_{i}|\leq 5.12\\ 5.12/\theta_{i}&\text{otherwise}\\ \end{cases}$ (6)
@@ 6.2 Domain Details
For DO experiments, we set the objective function $f(\bm{\theta})=1$. Following prior work (Fontaine and Nikolaidis,, 2021, 2023), we let the solution dimension be $n=100$.
Arm Repertoire. The goal of Arm Repertoire (Cully and Demiris,, 2017; Vassiliades and Mouret,, 2018) is to search for a diverse collection of arm positions for a planar robotic arm with $n$ revolving joints. In this domain, ${\bm{\theta}}\in[-\pi,\pi]^{n}$ represents the angles of the $n$ joints, and ${\bm{\phi}}({\bm{\theta}})$ computes the $(x,y)$ position of the arm’s end-effector using forward kinematics. While all other domains in this paper have a maximum of $100\%$ coverage, Arm Repertoire has a maximum coverage of $80.24\%$ when using a $100\times 100$ grid archive (Fontaine and Nikolaidis,, 2023), since the arm can only move in a circle of radius $n$. Similar to LP, we set $f({\bm{\theta}})=1$ and $n=100$.
@@
(b) The layout in the deceptive maze domain. The red dot indicates the start, and the green dot indicates the goal. While there is a goal, DO seeks to find agents that reach every position in the maze. Adapted from prior work (Lehman and Stanley,, 2011).
Figure 2:
Deceptive Maze. Deceptive Maze is a DO domain that challenges the algorithm to discover a diverse set of final positions for robots navigating a maze (1(b)(Lehman and Stanley,, 2011). In this domain, ${\bm{\theta}}$ parameterizes the robot’s neural network controller. ${\bm{\phi}}({\bm{\theta}})$ is the final position of the robot after evaluating its path in the maze. As a DO domain, this has no objective function. In our experiments, the neural network is a MLP with $n=66$ parameters.
Multi-feature Linear Projection (Multi-feature LP). We introduce a generalized version of LP that scales to $m$-dimensional feature spaces. The feature function ${\bm{\phi}}:\mathbb{R}^{n}\to\left[-5.12\cdot\frac{n}{m},5.12\cdot\frac{n}{m}\right]^{m}$ is defined as
${\bm{\phi}}({\bm{\theta}})=\left(\sum_{i=\frac{jn}{m}+1}^{\frac{(j+1)n}{m}}\mathrm{clip}(\theta_{i}):j\in\{0,\dots,m\}\right)$ (7)
where we assume that $n$ is divisible by $m$. When $m=2$, this is equivalent to LP definition in Eq. 5. Our experiments use $n=100$ and $m=10$.
6.3 Analysis

### v2

@@ 4. Density Descent Search
The second variant, DDS-CNF, estimates the density in feature space by learning a CNF between the standard normal distribution $\mathcal{N}(\bm{0},{\bm{I}})$ and the observed feature space distribution. Similar to DDS-KDE, the feature space distribution is represented by the buffer ${\mathcal{B}}$. We compute a density estimate at any location in the feature space by integrating an ordinary differential equation (Lipman et al., 2023). Applying techniques from prior work (Lipman et al., 2023), we represent the CNF with a neural network, and on each iteration, we finetune the network on the new distribution of features contained in the buffer.
KDE and CNF differ in how easily their smoothness can be controlled. On one hand, the shape of the KDE can be controlled with the bandwidth hyperparameter. Higher bandwidth leads to a smoother density estimate but conceals modes of the true density function, as illustrated in 2(a). On the other hand, while CNF does not require selecting a bandwidth hyperparameter, the smoothness of the density estimation cannot be easily controlled, which undermines the performance of the algorithm (Sec. 6.3). We discuss the impacts that the smoothness of the density estimate has on DDS in Sec. 5.
Feature Buffer. To provide the basis for density estimates in feature space, DDS maintains a buffer ${\mathcal{B}}$ that stores the features of sampled solutions (Algorithm 1). This buffer represents the observed distribution of the feature space and is updated every time new solutions and features are discovered. In theory, the buffer can have infinite capacity, storing every feature ever encountered. In practice, the buffer can only retain a finite number of features due to computation and memory limitations. To decide which features to retain in the buffer, we manage the buffer with an optimal reservoir sampling algorithm (Li, 1994). This algorithm updates the buffer with online samples that accurately represent the distribution of features discovered so far by DDS (Algorithm 1).
Optimizer. DDS guides an adaptive optimizer to discover solutions in low-density regions of the feature space. Examples of such optimizers include xNES (Glasmachers et al., 2010) and Adam (Kingma and Ba, 2015). We select CMA-ES as the underlying optimizer due to its reputation as a state-of-the-art optimizer (Hansen, 2016) and its high performance in existing QD algorithms (Fontaine et al., 2020; Fontaine and Nikolaidis, 2021, 2023).
Passive Archive. Following prior work (Pugh et al., 2015), to track how much of the feature space has been explored, DDS inserts all discovered solutions and features into a passive MAP-Elites-style (Sec. 3) archive $\mathcal{A}$. While DDS itself only uses the archive to record solutions and features, the archive is useful when computing metrics in experiments (Sec. 6.1).
@@
1 DDS ($\bm{\phi}(\cdot)$, $density$, $b,N,\lambda,{\bm{\mu}}_{0},\sigma$)
@@
2
3      Initialize CMA-ES search point ${\bm{\mu}}:={\bm{\mu}}_{0}$, search direction ${\bm{\Sigma}}=\sigma{\bm{I}}$, and internal parameters ${\bm{p}}$
4      Initialize empty archive $\mathcal{A}$
5      Initialize empty buffer ${\mathcal{B}}$ of size $b$
6      for $iter\leftarrow 1$ to $N$ do
7             Update buffer ${\mathcal{B}}$ with ${\bm{\phi}}_{1..\lambda}$ via reservoir sampling
@@ 5      Initialize empty buffer ${\mathcal{B}}$ of size $b$
9                   ${\bm{\theta}}_{i}\sim\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$
10                  ${\bm{\phi}}_{i}\leftarrow{\bm{\phi}}({\bm{\theta}}_{i})$
11                  $D_{i}\leftarrow density({\bm{\phi}}_{i})$
12                  Add $({\bm{\theta}}_{i},{\bm{\phi}}_{i})$ to archive $\mathcal{A}$
13            Update $density$ with the new buffer ${\mathcal{B}}$
14            Rank ${\bm{\theta}}_{i}$ in ascending order by $D_{i}$
15            Adapt CMA-ES parameters ${\bm{\mu}},{\bm{\Sigma}},{\bm{p}}$ based on density ranking $D_{i}$
@@
5. Connection between Novelty Search and Kernel Density Estimates
@@
Stability Under Non-stationarity. DO algorithms such as NS, DDS, and CMA-MAE gradually learn a non-stationary density representation (e.g., KDE or histogram) as they explore the feature space. However, drastic changes in the representation present significant challenges for adaptive optimizers, such as Adam (Kingma and Ba, 2015) and CMA-ES. If the density representation changes drastically every iteration, it would be impossible for adaptive optimizers to properly adapt its parameters to optimize the density function.
To characterize the extent of change in the density estimate, we appeal to the notion of uniform stability (Hardt et al., 2016), defined as follows:
@@
Let a function $F({\bm{x}};{\mathcal{B}}):\mathbb{R}^{d}\to\mathbb{R}$ be parameterized by a set ${\mathcal{B}}\subseteq\mathbb{R}^{d}$. We say that $F$ is $\epsilon$-uniformly stable if for all ${\mathcal{B}},{\mathcal{B}}^{\prime}\subseteq\mathbb{R}^{d}$, where ${\mathcal{B}}$ and ${\mathcal{B}}^{\prime}$ differ by at most one element, we have
(3) $\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left|F({\bm{x}};{\mathcal{B}})-F({\bm{x}};{\mathcal{B}}^{\prime})\right|\leq\epsilon$
We note that uniform stability has close connections to influence functions in statistics, which measure how much an estimator deviates from the ground truth when given a subset of the data (Cook and Weisberg, 1980). In this interpretation, uniform stability is an upper-bound of the (empirical) influence function.
We prove that, for KDE, the higher the size of its feature buffer and bandwidth, the stronger the uniform stability guarantee.
@@
Higher bandwidth makes the function more stationary and thus more suited for adaptive optimizers. However, higher bandwidth also leads to over-smoothing of the density estimate, which conceals modes of the true density function (Chen, 2017) (2(a)). Thus, selecting an optimal bandwidth for DDS requires a fine balance between the accuracy and uniform stability of the KDE.
@@
Therefore, for unbounded feature spaces, the uniform stability of novelty score is also unbounded, and for bounded feature spaces, novelty score has stronger uniform stability guarantees when $k$ is larger.
When the feature space is bounded, as in our experiments, KDE has a stronger stability guarantee than novelty score for bandwidth $h\geq\frac{k}{|{\mathcal{B}}|}$. Following this insight, we select a bandwidth satisfying this inequality with the empirical experiments in Appendix B. Our theoretical results are corroborated by our experiments in Sec. 6, which demonstrate that DDS-KDE outperforms NS in all domains on all metrics.
@@
Let ${\mathcal{B}}\subseteq\mathbb{R}^{m}$ be a set of features. Consider the rankings $\pi_{\mathrm{NS}}$ and $\pi_{\mathrm{KDE}}$ on another set of features $\{{\bm{\phi}}_{1},\dots,{\bm{\phi}}_{m}\}\subseteq\mathbb{R}^{m}$ where ${\bm{\phi}}_{i}={\bm{\phi}}({\bm{\theta}}_{i})$. We define the rankings as follows: $\pi_{\mathrm{NS}}(i)\geq\pi_{\mathrm{NS}}(j)\iff\rho({\bm{\phi}}_{i};{\mathcal{B}})\geq\rho({\bm{\phi}}_{j};{\mathcal{B}})$ and $\pi_{\mathrm{KDE}}(i)\geq\pi_{\mathrm{KDE}}(j)\iff\hat{D}({\bm{\phi}}_{i};{\mathcal{B}})\leq\hat{D}({\bm{\phi}}_{j};{\mathcal{B}})$. We show that $\pi_{NS}=\pi_{KDE}$, when NS has $k=\infty$ (or, equivalently, $k=|{\mathcal{B}}|$) and KDE has triangular kernel $K(u)=1-|u|$ with support over the entire feature space $S$.
@@
6. Experiments
We compare the performance of DDS-KDE and DDS-CNF with the QD algorithms MAP-Elites (line), CMA-ME, and CMA-MAE, and with NS using CMA-ES as the underlying optimizer. MAP-Elites (line) refers to the implementation of MAP-Elites with the Iso+LineDD operator (Vassiliades and Mouret, 2018). All algorithms are implemented with pyribs (Tjanaka et al., 2023).
Our experiments include canonical benchmark domains from QD and DO: Linear Projection (Fontaine et al., 2020), Arm Repertoire (Vassiliades and Mouret, 2018), and Deceptive Maze (Lehman and Stanley, 2011). As discussed in Sec. 2, we convert QD domains into DO domains by setting the objective to be constant, effectively removing the importance of solution quality. For Deceptive Maze, we use the implementation in Kheperax (Grillotti and Cully, 2023). Furthermore, to address the challenges posed by high-dimensional feature spaces, we introduce a new domain, Multi-feature Linear Projection, that generalizes Linear Projection to high-dimensional feature spaces.
@@ 6.1. Experimental Design
Dependent Variables. We compute the archive coverage as the number of occupied cells in the archive divided by the total number of cells. To compare the coverage of tessellation-based algorithms (CMA-MAE, CMA-ME, and MAP-Elites) with that of non-tessellation-based algorithms (DDS-KDE, DDS-CNF, and NS), we track the coverage of all algorithms on a passive archive $\mathcal{A}$ tessellated by a $100\times 100$ grid. For the high-dimensional feature space experiment, we track coverage with a centroidal Voronoi tessellation with 10,000 cells (Vassiliades et al., 2018).
Following prior work (Gomes et al., 2015), we also assess the ability of each algorithm to uniformly explore the feature space. Thus, we measure the cross-entropy between a uniform distribution and the distribution of sampled features. Let $N_{e}$ be the total times throughout the entire search that solutions were discovered in cell $e\in\{1,\dots,l\}$ in the passive archive, and $N_{\mathrm{total}}=\sum_{e=1}^{l}N_{e}$. The cross-entropy score is defined as:
(4) $\mathrm{CE}=-\sum_{e=1}^{l}\frac{1}{l}\log\left(\frac{N_{e}}{N_{\mathrm{total}}}\right)$
$\mathrm{CE}$ achieves its minimum value when $N_{e}$ is uniformly distributed across all cells $e\in\{1,\dots,l\}$. For all passive archives in our experiments, this minimum is 9.21 for $l=$10,000$$ cells.
6.2. Domain Details
Linear Projection (LP). LP is a QD domain where ${\bm{\phi}}$ induces high distortion by mapping an $n$-dimensional solution space to a 2D feature space (Fontaine et al., 2020). A harsh penalty is applied outside the bounds of the feature space to hinder exploration near the bounds. The feature function ${\bm{\phi}}:\mathbb{R}^{n}\to\left[-5.12\cdot\frac{n}{2},5.12\cdot\frac{n}{2}\right]^{2}$ is defined as:
(5) ${\bm{\phi}}({\bm{\theta}})$ $=\left(\sum_{i=1}^{\frac{n}{2}}\mathrm{clip}(\theta_{i}),\sum_{i=\frac{n}{2}+1}^{n}\mathrm{clip}(\theta_{i})\right)$
(6) $\mathrm{clip}(\theta_{i})$ =$=\begin{cases}\theta_{i}&\text{if}\;|\theta_{i}|\leq 5.12\\ 5.12/\theta_{i}&\text{otherwise}\\ \end{cases}$
@@ 6.2. Domain Details
For DO experiments, we set the objective function $f(\bm{\theta})=1$. Following prior work (Fontaine and Nikolaidis, 2021, 2023), we let the solution dimension be $n=100$.
Arm Repertoire. The goal of Arm Repertoire (Cully and Demiris, 2017; Vassiliades and Mouret, 2018) is to search for a diverse collection of arm positions for a planar robotic arm with $n$ revolving joints. In this domain, ${\bm{\theta}}\in[-\pi,\pi]^{n}$ represents the angles of the $n$ joints, and ${\bm{\phi}}({\bm{\theta}})$ computes the $(x,y)$ position of the arm’s end-effector using forward kinematics. While all other domains in this paper have a maximum of $100\%$ coverage, Arm Repertoire has a maximum coverage of $80.24\%$ when using a $100\times 100$ grid archive (Fontaine and Nikolaidis, 2023), since the arm can only move in a circle of radius $n$. Similar to LP, we set $f({\bm{\theta}})=1$ and $n=100$.
@@
(b) The layout in the deceptive maze domain. The red dot indicates the start, and the green dot indicates the goal. While there is a goal, DO seeks to find agents that reach every position in the maze. Adapted from prior work (Lehman and Stanley, 2011).
Figure 2.
Deceptive Maze. Deceptive Maze is a DO domain that challenges the algorithm to discover a diverse set of final positions for robots navigating a maze (2(b)(Lehman and Stanley, 2011). In this domain, ${\bm{\theta}}$ parameterizes the robot’s neural network controller. ${\bm{\phi}}({\bm{\theta}})$ is the final position of the robot after evaluating its path in the maze. As a DO domain, this has no objective function. In our experiments, the neural network is a MLP with $n=66$ parameters.
Multi-feature Linear Projection (Multi-feature LP). We introduce a generalized version of LP that scales to $m$-dimensional feature spaces. Assuming that $n$ is divisible by $m$, let $r=\frac{n}{m}$. The feature function ${\bm{\phi}}:\mathbb{R}^{n}\to\left[-5.12\cdot r,5.12\cdot r\right]^{m}$ is defined as
(7) ${\bm{\phi}}({\bm{\theta}})=\left(\sum_{i=jr+1}^{(j+1)r}\mathrm{clip}(\theta_{i}):j\in\{0,\dots,m-1\}\right)$
When $m=2$, this is equivalent to LP definition in Eq. 5. Our experiments use $n=100$ and $m=10$.
6.3. Analysis